# Dynamical Edge Modes in $p$-form Gauge Theories

### v1

@@ 4.3.1 Log-coefficients in even $D$
$D$ $p$ 00 $1$ $2$ $3$ $4$ $5$ $6$ $7$ $8$
$4$ 00 $-\frac{1}{3}$ $2$
$6$ 00 $-\frac{29}{90}$ $\frac{31}{45}$ $-\frac{209}{90}$ $4$
$8$ 00 $-\frac{1139}{3780}$ $\frac{1271}{1890}$ $-\frac{221}{210}$ $\frac{5051}{1890}$ $-\frac{16259}{3780}$ $6$
$10$ 00 $-\frac{32377}{113400}$ $\frac{4021}{6300}$ $-\frac{2603}{2520}$ $\frac{8051}{5670}$ $-\frac{7643}{2520}$ $\frac{29221}{6300}$ $-\frac{712777}{113400}$ $8$
@@
Except for $p=0$ and $p=D-2$, table 4.3 is the same as table 4.1 after shifting $p\to p-1$ and $D\to D-2$ and taking the negatives. One can readily check that $\log\mathcal{Z}^{p}_{\rm PI,log}\left[S^{D}\right]=\log\mathcal{Z}^{p}_{\rm bulk,log}\left[dS_{D}\right]+\log\mathcal{Z}^{p}_{\rm edge,log}\left[S^{D}\right]$.
@@
In tables 4.1 and 4.3, the coefficients for the $p$-form differ from those of the dual $(D-2-p)$-form by $(-)^{p}\left(D-2-2p\right)$. This “duality anomaly” in even $D$ was first observed in [75], where it was found that the trace anomalies for the 0- and 2-form in $D=4$ differ by the Euler number $\chi$ of the manifold. Later this was proved more generally in [76, 48].
@@
$=$ $\left(\frac{\mu^{\tilde{p}}\sqrt{2\pi R^{D}{\rm Vol}\left(S^{D}\right)}}{q\mu^{D-2-2p}}\right)^{(-)^{\tilde{p}}}\left(\det^{\prime}\frac{\Delta_{0}}{\mu^{2}}\right)^{-\frac{(-)^{\tilde{p}}}{2}}\prod_{k=1}^{\tilde{p}}\left(\det\frac{\Delta^{T}_{k}}{\mu^{2}}\right)^{-\frac{(-)^{\tilde{p}-k}}{2}}\;.$ (4.38)
@@
$\mathcal{Z}^{\tilde{p}}_{\rm PI}\left[S^{D}\right]=\left(\frac{\mu^{\tilde{p}}\sqrt{2\pi R^{D}{\rm Vol}\left(S^{D}\right)}}{\tilde{q}}\right)^{(-)^{\tilde{p}}}\left(\det^{\prime}\frac{\Delta_{0}}{\mu^{2}}\right)^{-\frac{(-)^{\tilde{p}}}{2}}\prod_{k=1}^{\tilde{p}}\left(\det\frac{\Delta^{T}_{k}}{\mu^{2}}\right)^{-\frac{(-)^{\tilde{p}-k}}{2}}\;.$ (4.39)
@@
$\log\frac{\mathcal{Z}^{p}_{\rm PI}\left[S^{D}\right]}{\mathcal{Z}^{\tilde{p}}_{\rm PI}\left[S^{D}\right]}=$ $(-)^{\tilde{p}+1}2\log\left(\mu^{\frac{D-2}{2}-p}\sqrt{\frac{q}{\tilde{q}}}\right)$
@@ 4.3.2 UV-finite parts in odd $D$
00 $-\frac{\zeta(3)}{4\pi^{2}}$ $\log\frac{2\pi\sqrt{R}}{q}$ $\log\frac{2\pi\sqrt{R}}{q}-\frac{\zeta(3)}{4\pi^{2}}$
$1$ $-\frac{\zeta(3)}{4\pi^{2}}$ $\log q\sqrt{R}$ $\log q\sqrt{R}-\frac{\zeta(3)}{4\pi^{2}}$
@@
0 $-\frac{23\zeta(3)}{48\pi^{2}}+\frac{\zeta(5)}{16\pi^{4}}$ $\log\frac{2\pi^{3/2}R^{3/2}}{q}$ $\log\frac{2\pi^{3/2}R^{3/2}}{q}-\frac{23\zeta(3)}{48\pi^{2}}+\frac{\zeta(5)}{16\pi^{4}}$
@@
3 $-\frac{23\zeta(3)}{48\pi^{2}}+\frac{\zeta(5)}{16\pi^{4}}$ $\log\left(qR^{3/2}\pi^{1/2}\right)$ $\log\left(qR^{3/2}\pi^{1/2}\right)-\frac{23\zeta(3)}{48\pi^{2}}+\frac{\zeta(5)}{16\pi^{4}}$
@@
0 $-\frac{949\,\zeta(3)}{1440\,\pi^{2}}+\frac{13\,\zeta(5)}{48\,\pi^{4}}-\frac{\zeta(7)}{64\,\pi^{6}}$ $\log\frac{\sqrt{2}\,\pi^{2}\,R^{5/2}}{q}$ $\log\frac{\sqrt{2}\,\pi^{2}\,R^{5/2}}{q}-\frac{949\,\zeta(3)}{1440\,\pi^{2}}+\frac{13\,\zeta(5)}{48\,\pi^{4}}-\frac{\zeta(7)}{64\,\pi^{6}}$
@@
2 $-\frac{\zeta(3)}{36\,\pi^{2}}-\frac{5\,\zeta(5)}{48\,\pi^{4}}-\frac{5\,\zeta(7)}{32\,\pi^{6}}$ $\log\frac{2\,\pi\,R^{1/2}}{q}-\frac{5\,\zeta(3)}{16\,\pi^{2}}-\frac{3\,\zeta(5)}{16\,\pi^{4}}$ $\log\frac{2\,\pi\,R^{1/2}}{q}-\frac{49\,\zeta(3)}{144\,\pi^{2}}-\frac{7\,\zeta(5)}{24\,\pi^{4}}-\frac{5\,\zeta(7)}{32\,\pi^{6}}$
3 $-\frac{\zeta(3)}{36\,\pi^{2}}-\frac{5\,\zeta(5)}{48\,\pi^{4}}-\frac{5\,\zeta(7)}{32\,\pi^{6}}$ $\log q\sqrt{R}-\frac{5\,\zeta(3)}{16\,\pi^{2}}-\frac{3\,\zeta(5)}{16\,\pi^{4}}$ $\log q\sqrt{R}-\frac{49\,\zeta(3)}{144\,\pi^{2}}-\frac{7\,\zeta(5)}{24\,\pi^{4}}-\frac{5\,\zeta(7)}{32\,\pi^{6}}$
@@
5 $-\frac{949\,\zeta(3)}{1440\,\pi^{2}}+\frac{13\,\zeta(5)}{48\,\pi^{4}}-\frac{\zeta(7)}{64\,\pi^{6}}$ $\log\frac{q\,\pi\,R^{5/2}}{\sqrt{2}}$ $\log\frac{q\,\pi\,R^{5/2}}{\sqrt{2}}-\frac{949\,\zeta(3)}{1440\,\pi^{2}}+\frac{13\,\zeta(5)}{48\,\pi^{4}}-\frac{\zeta(7)}{64\,\pi^{6}}$
@@ 4.4 Entanglement entropy for even spheres and the conformal case
We will (partially) prove this for any $p$ and even $D$ below. Granting this, the bulk-edge split (4.28) of the brick-wall regulated DEM partition function implies a bulk-edge split for the entanglement entropy as well. This is in accordance with earlier calculations of entanglement entropy [48, 49, 50], which suggested the existence of edge contributions from a $(p-1)$-form residing on the codimension-2 entangling surface. Our findings corroborate these earlier observations and offer a dynamical framework to understand them as large gauge transformations and their symplectic conjugates, captured by the DEM boundary condition (3.3).
@@
For example, the $k=0$ term of (4.26) takes the form (4.49) with $\Delta=0$. For these values of $\Delta$, (4.52) vanishes. Note that the $-(-)^{p}$ term in (4.26) does not alter the coefficient of $\frac{1}{t}$ in the small-$t$ expansion. This establishes the relation (4.45) (up to a proof of (4.51)).
@@ 4.5.1 Massive $p$-forms
We first study a massive $p$-form on the Euclidean manifold $\mathcal{M}$ considered in section 4.2, namely one obtained from the static Lorentzian manifold $M$ with metric (3.37) by a Wick rotation $t\to-i\tau$ with $\tau\sim\tau+2\pi$. Following section 4.2, we once again cut out a hole of proper radius $\varepsilon$ encircling the origin, resulting in a manifold $\mathcal{M}_{\varepsilon}$. Following [10], we would like to compute the partition function [^17] Classically, a massive $p$-form is dual to a massive $(D-1-p)$-form. See for instance [81, 82] for discussions on their quantum equivalence.
@@
where ^$\hat{\nabla}_{i}$ is the covariant derivative with respect to the metric $g_{ij}$ on $\Sigma$, and we have defined
@@
This $(p-1)$-form turns out to play a highly analogous role to the edge mode $\beta$ in the $p$-form gauge theory in section 3. Imposing the equation of motion (4.58), it obeys
^$\hat{\nabla}^{j}(\sqrt{g^{\tau\tau}}\gamma_{ji_{1}\dots i_{p-2}})=0\;,\qquad m^{2}\gamma_{i_{1}\dots i_{p-1}}=\frac{1}{\sqrt{g^{\tau\tau}}}\hat{\nabla}^{j}\left(\sqrt{g^{\tau\tau}}p!\hat{\nabla}_{[j}\gamma_{i_{1}\dots i_{p-1}]}\right)\;.$ (4.65)
@@
$\mathcal{D}E_{\perp}=\mathcal{D}E_{\perp}^{T}\mathcal{D}(\bar{d}\chi)=\mathcal{D}E_{\perp}^{T}\mathcal{D}\chi\det^{\prime}(\Delta^{T,\partial\Sigma_{\varepsilon}}_{p-2})^{1/2}\;.$ (4.76)
@@
$Z^{p,m^{2}}_{\rm edge}[\mathcal{M}_{\varepsilon}]$ $=\det^{\prime}\left(\Delta^{T,\partial\Sigma_{\varepsilon}}_{p-2}\right)^{1/2}\int\mathcal{D}E_{\perp}^{T}\int\mathcal{D}\chi\,e^{-S[E_{\perp}^{T}]-S[\bar{d}\chi]}$ (4.77)
$=\det^{\prime}\left(\Delta^{T,\partial\Sigma_{\varepsilon}}_{p-2}\right)^{1/2}\det(\frac{\log(\varepsilon^{-1})}{2\pi\mu^{2}}(\Delta^{T,\partial\Sigma_{\varepsilon}}_{p-1}+m^{2}))^{1/2}\det^{\prime}\left(\frac{m^{2}\log(\varepsilon^{-1})}{2\pi\mu^{2}}\Delta^{T,\partial\Sigma_{\varepsilon}}_{p-2}\right)^{-1/2}\;.$
@@
where $\mathcal{A}$, $\mathcal{A}^{\prime}$ are some in principle computable anomalies present only when $D$ is even. Up to the constant factors out front, we have the (reciprocal) partition function of a massive $(p-1)$-form on $\partial\Sigma_{\varepsilon}$. To summarize, the object
@@
defined in terms of the Harish-Chandra character $\chi(t)$ for a $p$-form with mass $m^{2}R^{2}=\left(\Delta-p\right)\left(\bar{\Delta}-p\right)=\left(\Delta-p\right)\left(D-1-\Delta-p\right)$. The edge partition function $\mathcal{Z}^{p,m^{2}}_{\rm edge}$ is the reciprocal of a $(p-1)$-form partition function with mass $m^{2}$ on $S^{D-2}$:

### v2

@@ 66 $\frac{1139}{3780}$ $\frac{331}{945}$ $\frac{229}{630}$ $\frac{331}{945}$ $\frac{1139}{3780}$
$D$ pp 0 11 22 33 44 55 66 77 88
44 0 $\frac{1}{3}$ 22
66 0 $\frac{29}{90}$ $\frac{31}{45}$ $\frac{209}{90}$ 44
88 0 $\frac{1139}{3780}$ $\frac{1271}{1890}$ $\frac{221}{210}$ $\frac{5051}{1890}$ $\frac{16259}{3780}$ 66
1010 0 $\frac{32377}{113400}$ $\frac{4021}{6300}$ $\frac{2603}{2520}$ $\frac{8051}{5670}$ $\frac{7643}{2520}$ $\frac{29221}{6300}$ $\frac{712777}{113400}$ 88
@@
Except for $p=0$ and p=D2p=D-2, table 4.3 is the same as table 4.1 after shifting $p\to p-1$ and $D\to D-2$ and taking the negatives. One can readily check that $\log\mathcal{Z}^{p}_{\rm PI,log}\left[S^{D}\right]=\log\mathcal{Z}^{p}_{\rm bulk,log}\left[dS_{D}\right]+\log\mathcal{Z}^{p}_{\rm edge,log}\left[S^{D}\right]$.
@@
In tables 4.1 and 4.3, the coefficients for the $p$-form differ from those of the dual (D2p)(D-2-p)-form by $(-)^{p}\left(D-2-2p\right)$. This “duality anomaly” in even $D$ was first observed in [88], where it was found that the trace anomalies for the 0- and 2-form in $D=4$ differ by the Euler number $\chi$ of the manifold. Later this was proved more generally in [89, 50].
@@
$=$ $\left(\frac{\mu^{\tilde{p}}\sqrt{2\pi R^{D}{\rm Vol}\left(S^{D}\right)}}{q\mu^{D-2-2p}}\right)^{(-)^{\tilde{p}}}\left(\det\hskip 0.0pt^{\prime}\frac{\Delta_{0}}{\mu^{2}}\right)^{-\frac{(-)^{\tilde{p}}}{2}}\prod_{k=1}^{\tilde{p}}\left(\det\frac{\Delta^{T}_{k}}{\mu^{2}}\right)^{-\frac{(-)^{\tilde{p}-k}}{2}}\;.$ (4.38)
@@
$\mathcal{Z}^{\tilde{p}}_{\rm PI}\left[S^{D}\right]=\left(\frac{\mu^{\tilde{p}}\sqrt{2\pi R^{D}{\rm Vol}\left(S^{D}\right)}}{\tilde{q}}\right)^{(-)^{\tilde{p}}}\left(\det\hskip 0.0pt^{\prime}\frac{\Delta_{0}}{\mu^{2}}\right)^{-\frac{(-)^{\tilde{p}}}{2}}\prod_{k=1}^{\tilde{p}}\left(\det\frac{\Delta^{T}_{k}}{\mu^{2}}\right)^{-\frac{(-)^{\tilde{p}-k}}{2}}\;.$ (4.39)
@@
[…]
@@ 4.3.2 UV-finite parts in odd $D$
0 $\frac{\zeta(3)}{4\pi^{2}}$ $\log\frac{2\pi\sqrt{R}}{q}$ $\log\frac{2\pi\sqrt{R}}{q}-\frac{\zeta(3)}{4\pi^{2}}$
11 $\frac{\zeta(3)}{4\pi^{2}}$ $\log q\sqrt{R}$ $\log q\sqrt{R}-\frac{\zeta(3)}{4\pi^{2}}$
@@
0 $\frac{23\zeta(3)}{48\pi^{2}}+\frac{\zeta(5)}{16\pi^{4}}$ $\log\frac{2\pi^{3/2}R^{3/2}}{q}$ $\log\frac{2\pi^{3/2}R^{3/2}}{q}-\frac{23\zeta(3)}{48\pi^{2}}+\frac{\zeta(5)}{16\pi^{4}}$
@@
3 $\frac{23\zeta(3)}{48\pi^{2}}+\frac{\zeta(5)}{16\pi^{4}}$ $\log\left(qR^{3/2}\pi^{1/2}\right)$ $\log\left(qR^{3/2}\pi^{1/2}\right)-\frac{23\zeta(3)}{48\pi^{2}}+\frac{\zeta(5)}{16\pi^{4}}$
@@
0 $\frac{949\,\zeta(3)}{1440\,\pi^{2}}+\frac{13\,\zeta(5)}{48\,\pi^{4}}-\frac{\zeta(7)}{64\,\pi^{6}}$ $\log\frac{\sqrt{2}\,\pi^{2}\,R^{5/2}}{q}$ $\log\frac{\sqrt{2}\,\pi^{2}\,R^{5/2}}{q}-\frac{949\,\zeta(3)}{1440\,\pi^{2}}+\frac{13\,\zeta(5)}{48\,\pi^{4}}-\frac{\zeta(7)}{64\,\pi^{6}}$
@@
2 $\frac{\zeta(3)}{36\,\pi^{2}}-\frac{5\,\zeta(5)}{48\,\pi^{4}}-\frac{5\,\zeta(7)}{32\,\pi^{6}}$ $\log\frac{2\,\pi\,R^{1/2}}{q}-\frac{5\,\zeta(3)}{16\,\pi^{2}}-\frac{3\,\zeta(5)}{16\,\pi^{4}}$ $\log\frac{2\,\pi\,R^{1/2}}{q}-\frac{49\,\zeta(3)}{144\,\pi^{2}}-\frac{7\,\zeta(5)}{24\,\pi^{4}}-\frac{5\,\zeta(7)}{32\,\pi^{6}}$
3 $\frac{\zeta(3)}{36\,\pi^{2}}-\frac{5\,\zeta(5)}{48\,\pi^{4}}-\frac{5\,\zeta(7)}{32\,\pi^{6}}$ $\log q\sqrt{R}-\frac{5\,\zeta(3)}{16\,\pi^{2}}-\frac{3\,\zeta(5)}{16\,\pi^{4}}$ $\log q\sqrt{R}-\frac{49\,\zeta(3)}{144\,\pi^{2}}-\frac{7\,\zeta(5)}{24\,\pi^{4}}-\frac{5\,\zeta(7)}{32\,\pi^{6}}$
@@
5 $\frac{949\,\zeta(3)}{1440\,\pi^{2}}+\frac{13\,\zeta(5)}{48\,\pi^{4}}-\frac{\zeta(7)}{64\,\pi^{6}}$ $\log\frac{q\,\pi\,R^{5/2}}{\sqrt{2}}$ $\log\frac{q\,\pi\,R^{5/2}}{\sqrt{2}}-\frac{949\,\zeta(3)}{1440\,\pi^{2}}+\frac{13\,\zeta(5)}{48\,\pi^{4}}-\frac{\zeta(7)}{64\,\pi^{6}}$
@@ 4.4 Entanglement entropy for even spheres and the conformal case
We will (partially) prove this for any $p$ and even $D$ below. Granting this, the bulk-edge split (4.28) of the brick-wall regulated DEM partition function implies a bulk-edge split for the entanglement entropy as well. This is in accordance with earlier calculations of entanglement entropy [50, 51, 52], which suggested the existence of edge contributions from a (p1)(p-1)-form residing on the codimension-2 entangling surface. Our findings corroborate these earlier observations and offer a dynamical framework to understand them as large gauge transformations and their symplectic conjugates, captured by the DEM boundary condition (3.3).
@@
For example, the $k=0$ term of (4.26) takes the form (4.49) with $\Delta=0$. For these values of $\Delta$, (4.52) vanishes. Note that the $(-)^{p}$ term in (4.26) does not alter the coefficient of $\frac{1}{t}$ in the small-$t$ expansion. This establishes the relation (4.45) (up to a proof of (4.51)).
@@ 4.5.1 Massive $p$-forms
We first study a massive $p$-form on the Euclidean manifold $\mathcal{M}$ considered in section 4.2, namely one obtained from the static Lorentzian manifold $M$ with metric (3.37) by a Wick rotation $t\to-i\tau$ with $\tau\sim\tau+2\pi$. Following section 4.2, we once again cut out a hole of proper radius $\varepsilon$ encircling the origin, resulting in a manifold $\mathcal{M}_{\varepsilon}$. Following [10], we would like to compute the partition function [^17] Classically, a massive $p$-form is dual to a massive (D1p)(D-1-p)-form. See for instance [94, 95] for discussions on their quantum equivalence.
@@
where $\hat{\nabla}_{i}$ is the covariant derivative with respect to the metric $g_{ij}$ on $\Sigma$, and we have defined
@@
This (p1)(p-1)-form turns out to play a highly analogous role to the edge mode $\beta$ in the $p$-form gauge theory in section 3. Imposing the equation of motion (4.58), it obeys
$\hat{\nabla}^{j}(\sqrt{g^{\tau\tau}}\gamma_{ji_{1}\dots i_{p-2}})=0\;,\qquad m^{2}\gamma_{i_{1}\dots i_{p-1}}=\frac{1}{\sqrt{g^{\tau\tau}}}\hat{\nabla}^{j}\left(\sqrt{g^{\tau\tau}}p!\hat{\nabla}_{[j}\gamma_{i_{1}\dots i_{p-1}]}\right)\;.$ (4.65)
@@
$\mathcal{D}E_{\perp}=\mathcal{D}E_{\perp}^{T}\mathcal{D}(\bar{d}\chi)=\mathcal{D}E_{\perp}^{T}\mathcal{D}\chi\det\hskip 0.0pt^{\prime}(\Delta^{T,\partial\Sigma_{\varepsilon}}_{p-2})^{1/2}\;.$ (4.76)
@@
$Z^{p,m^{2}}_{\rm edge}[\mathcal{M}_{\varepsilon}]$ $=\det\hskip 0.0pt^{\prime}\left(\Delta^{T,\partial\Sigma_{\varepsilon}}_{p-2}\right)^{1/2}\int\mathcal{D}E_{\perp}^{T}\int\mathcal{D}\chi\,e^{-S[E_{\perp}^{T}]-S[\bar{d}\chi]}$ (4.77)
$=\det\hskip 0.0pt^{\prime}\left(\Delta^{T,\partial\Sigma_{\varepsilon}}_{p-2}\right)^{1/2}\det(\frac{\log(\varepsilon^{-1})}{2\pi\mu^{2}}(\Delta^{T,\partial\Sigma_{\varepsilon}}_{p-1}+m^{2}))^{1/2}\det\hskip 0.0pt^{\prime}\left(\frac{m^{2}\log(\varepsilon^{-1})}{2\pi\mu^{2}}\Delta^{T,\partial\Sigma_{\varepsilon}}_{p-2}\right)^{-1/2}\;.$
@@
where $\mathcal{A}$, $\mathcal{A}^{\prime}$ are some in principle computable anomalies present only when $D$ is even. Up to the constant factors out front, we have the (reciprocal) partition function of a massive (p1)(p-1)-form on $\partial\Sigma_{\varepsilon}$. To summarize, the object
@@
defined in terms of the Harish-Chandra character $\chi(t)$ for a $p$-form with mass $m^{2}R^{2}=\left(\Delta-p\right)\left(\bar{\Delta}-p\right)=\left(\Delta-p\right)\left(D-1-\Delta-p\right)$. The edge partition function $\mathcal{Z}^{p,m^{2}}_{\rm edge}$ is the reciprocal of a (p1)(p-1)-form partition function with mass $m^{2}$ on $S^{D-2}$: